# Current validation practice undermines surgical AI development

### v1

@@ 3. Discussion
Our work provides the first comprehensive, expert-driven taxonomy of validation pitfalls in surgical AI, supported by empirical evidence and experimental results. By linking methodological flaws to surgical risks, our framework highlights the need for rigorous validation to ensure safe and effective AI deployment. Our findings demonstrate that common validation practices frequently ignore the temporal and hierarchical structure of surgical data, leading to overconfident or clinically irrelevant conclusions. Through a Delphi process with experts from surgery, machine learning, biostatistics, and regulatory affairs, we specifically identified and contextualized 18 critical pitfalls. This multi-stakeholder approach ensured that the collected pitfalls are both technically sound and clinically meaningful, making them actionable for a broad range of players, from algorithm developers to regulatory reviewers.
@@
The collected pitfalls span all stages of the validation process. For example, at the data level, geographical imbalance of surgical and endoscopic datasets (Extended Data Fig. 1b) shows how limited geographic diversity can restrict representativeness, as many public benchmarks have historically been acquired from a narrow range of regions. Recent initiatives, such as the Critical View of Safety (CVS) Challenges 2024 and 2025 [alapatt2025sages], represent encouraging steps toward broader global inclusion, yet most benchmark datasets still predominantly reflect surgical practice patterns from Western regions. At the metric level, misalignment between chosen metrics and clinical objectives remains common, as metrics tailored to surgeon-specific needs or temporal aspects are often lacking. Consequently, studies frequently rely on simple frame-based scores such as Accuracy or DSC, which fail to capture the temporal and hierarchical complexity of surgical workflows and can obscure clinically relevant weaknesses. At the aggregation level, naïve frame-wise aggregation across temporally dependent data can mislead performance estimates, concealing critical failure modes and even altering algorithm rankings.
@@
In addition to establishing this taxonomy, we collected empirical evidence confirming that the identified pitfalls are widespread across the surgical AI literature. Our systematic screening revealed frequent issues such as unstratified aggregation, lack of uncertainty reporting, unjustified and unsuitable metric selection, and poor documentation of validation procedures. Despite temporality being inherent to intraoperative surgical workflows and despite the existence of temporal metrics, these were rarely used in practice. Generally speaking, although surgical video AI tasks often involve complex, structured data such as multiple classes, temporal dependencies, and hierarchical structure, validation in common practice frequently remains limited to simple Accuracy metrics. As shown by our experiments, naïve frame-wise validation without consideration of temporal aspects or data hierarchy can mask critical failure modes and lead to a substantial underestimation of uncertainty. Together, these findings underscore the need for a paradigm shift in common surgical AI validation practice.
@@
Our work comes with several limitations. While our taxonomy targets surgical video analysis, its generalizability to other temporally structured domains (e.g., cardiology) remains to be explored. Methodologically, although our Delphi expert consortium included more than 90 international experts and covered a broad range of expertises, it may not have captured the full diversity of surgical subfields and regulatory perspectives. Participation also varied across Delphi rounds, raising the potential for biases in the weighting and selection of pitfalls. In addition, our systematic literature review only covered 46 articles. However, although the sample size was limited, MICCAI is the leading conference for medical image analysis and computer-assisted interventions and can thus be considered representative for the field. It should further be noted that our experiments only tackled selected pitfalls. While pitfalls such as data leakage have been demonstrated in the broader machine learning community, we explicitly focused on aspects that are especially critical in surgery: aggregation under temporal and hierarchical structures. Other pitfalls are illustrated in smaller real-data analyses in Extended Data Figs. 1- 9. Finally, our present work does not yet provide concrete solutions for each pitfall. However, developing concrete recommendations will be the focus of future work of the consortium.
@@
Several open research directions emerge from our presented pitfall taxonomy. At a conceptual level, future work should move beyond surrogate metrics toward validation that reflects clinical benefit and patient outcomes. This includes defining what constitutes sufficient performance within specific clinical contexts and establishing comparability across heterogeneous tasks. Methodologically, clear validation phases for SDS systems that integrate governance and stakeholder input, as well as standardized reporting, need to be defined and appropriately integrated into clinical trial design. Ensuring robustness through post-deployment monitoring, addressing catastrophic-failure risk and safety considerations, and enabling effective human-AI collaboration in the OR will be equally crucial. From a technical and adoption perspective, progress depends on harmonizing label ontologies and annotation protocols across datasets, facilitating validation for multimodal data while protecting patient privacy, and assessing behavioral consistency of AI models across samples and software versions to ensure stability after updates or retraining. Finally, embedding clinician priorities, workflow impact, and real-time safety mechanisms should become integral validation goals.
@@
In summary, while Metrics Reloaded [maier2024metrics] provided metric recommendations for image-based validation, our new framework extends this foundation to pitfalls stemming from the temporal and hierarchical complexity of surgical video analysis. We envision this work as a catalyst for improved validation practice and future benchmarking efforts. By raising awareness of widespread pitfalls, we aim to inspire a paradigm shift toward more robust, interpretable, and clinically grounded validation pipelines. By systematically mapping validation pitfalls to their consequences, this work offers a structured foundation for integrating validation quality criteria into clinical trial design, regulatory review, and publication guidelines. Concretely, our findings may inform the development or refinement of reporting standards for medical video analysis, such as TRIPOD-AI [collins2021protocol], DECIDE-AI [vasey2022reporting], or future domain-specific extensions. Going forward, the consortium will focus on translating these pitfalls into surgery-specific metric and aggregation recommendations, further advancing the reliability and clinical readiness of surgical AI models.
@@ 4. Methods
4.1. Identification of validation pitfalls through a multi-stage Delphi process and complementary searches
@@
The pitfalls presented in this work were derived through a combination of approaches, centered on a multi-stage, consensus-driven Delphi process conducted by an international, multidisciplinary panel of experts. A Delphi process is a structured consensus-building approach in which experts provide input individually – typically through questionnaires – followed by rounds of controlled feedback and refinement [brown1968delphi]. This methodology is widely recognized in medicine as a way to establish best practices in areas in which the available evidence is limited, inconsistent, or missing [nasa2021delphi].
@@
Our Delphi panel initially included 60 international experts from the SDS initiative. To broaden the diversity of the expertise, the consortium was gradually expanded to 91 members across 68 institutions, reflecting both technical and clinical backgrounds. The expert panel was composed of 31% clinical, 74% technical, and 5% shared expertise. 12% of experts were from industry. The majority of experts were affiliated in Europe (69%; mostly Germany (35%) and United Kingdom (12%)) and North America (25%; mostly United States of America (20%)), followed by Asia (7%) and Africa (2%).
@@
This initiative started in March 2023 with an initial scoping survey to identify the most critical problems in validating surgical AI and to capture use cases lacking suitable metrics or showing discrepancies between metrics and the clinical needs (participation rate: 33%). Building on the survey findings, we held an in-person kickoff workshop at the SAGES annual meeting in Montréal, Canada (41% in-person participation). This workshop refined the project scope and set priorities, laying the groundwork for the following Delphi rounds. Based on the workshop discussions, participants agreed to focus the initiative on surgical video understanding, reflecting shared priorities across clinical and technical stakeholders. Following the workshop, the core team performed targeted literature searches to compile candidate pitfalls. In parallel, a joint retreat involving members from three research groups at the German Cancer Research Center (DFKZ), National Center for Tumor Diseases (NCT), and University College London (UCL) provided an additional forum to critically discuss and refine preliminary pitfalls based on practical experience and interdisciplinary perspectives.
@@
In addition, the core team performed a literature review to identify additional pitfalls. Concretely, we explored two databases, namely PubMed and GoogleScholar, as well as a general Google search, and used the following search string: (”surgical data science” OR ”surgical artificial intelligence” OR ”surgical AI” OR ”surgical scene understanding” OR ”surgical video analysis”) AND (”validation” OR ”evaluation” OR ”metric”) AND (”pitfall” OR ”limitation” OR ”caveat” OR ”drawback” OR ”shortcoming” OR ”weakness” OR ”flaw” OR ”disadvantage”). PubMed returned no relevant results, while Google Scholar yielded 704 hits and a general Google search 94 results (30 non-peer-reviewed). Although these searches yielded several appropriate items, no comprehensive or structured overview of validation pitfalls was identified.
@@
In total, we conducted four Delphi rounds (participation rates: 64%, 70%, 48%, 53%). Round 1 confirmed the overall project scope, while round 2 refined the pitfall categorization and pitfall list, identified missing pitfalls, and collected supporting references for the evidence of pitfalls. Round 3 focused on linking pitfalls to consequences and risks, and round 4 sought final consensus on the pitfall catalog and pitfall categorization (agreement: 98%), as well as optional feedback on figures, experiments (agreement 90%), and open research questions.
@@ 4.2. Systematic review for prevalence of pitfalls
From all MICCAI 2023 papers (n = 730), we identified all papers related to SDS (n = 51). From those, five articles were excluded because they did not deal with deep learning-related methods, therefore, several questions did not apply (n = 46). Each paper was screened by two independent screeners, with a total of twelve screeners. Afterwards, a third senior screener compared results and resolved conflicts. In total, three senior screeners joined this last step, with the papers divided among them. The screening covered more general aspects such as task or surgery type, but specifically focused on evidence for the identified pitfalls. In line with the Delphi process and after identifying the final list of pitfalls, a follow-up screening, following the same process, was conducted to ensure evidence for all pitfalls.
@@ 4.3. Experimental Design
To ensure the general relevance of our findings, we based our experiments on two of the most widely used datasets in surgical video analysis [carstens2025artificial] that have been used in international challenges, are highly cited in the field, and cover two key tasks in surgical video analysis: instrument segmentation and action recognition.
The RobustMIS challenge 2019 [ross2021comparative] consists of videos from 30 laparoscopic colorectal surgeries, covering three surgery types, namely rectal resection, proctocolectomy, and sigmoid resection, with 10 videos, i.e., patients, per surgery type. For the challenge, the data from rectal resection and proctocolectomy surgeries were used for training and internal testing (stages 1 and 2), while sigmoid resection cases (stage 3) were reserved for assessing generalization to an unseen surgery type. For our experiments, we restricted the analysis to data from stage 3 (sigmoid resection). The challenge consisted of three tasks: binary segmentation, multi-instance detection, and multi-instance segmentation. For our experiments, we leveraged results for both segmentation tasks. For binary segmentation, a total of ten algorithms participated, for multi-instance segmentation, seven algorithms participated. The challenge metrics included the DSC, which measures the overlap between prediction and reference, and NSD, which assesses boundary accuracy [maier2024metrics] for binary segmentation and their multi-instance variants (MI_DSC and MI_NSD) for the multi-instance segmentation task. We had access to the frame-level metric scores submitted by all participating teams, which allowed us to analyze the impact of validation choices across a diverse set of real-world algorithms. All results were used in anonymized form to ensure confidentiality.
In addition to the segmentation masks, we utilized additional structured meta-annotations for the RobustMIS data, describing the presence of common visual artifacts or image characteristics [ross2023beyond]. These meta annotations indicate whether specific challenges such as blood, smoke, or motion blur are present per frame and instrument.
We further used the data and task setup of the CholecTriplet challenge [nwoye2023cholectriplet2021], which is based on 45 laparoscopic cholecystectomy videos (CholecT45). The task involves recognition of surgical action triplets, with annotations for 100 triplet classes, each defined by a combination of instrument, verb, and target. For our experiments, we used a Swin-Base Transformer trained using multitask learning, incorporating information on the instrument, verb, and target and soft-labels generated using a multi-teacher approach [yamlahi2023self]. Model validation was performed using 5-fold cross-validation, following the official CholecT45 setup.
@@
This experiment investigated how ignoring data dependencies in temporally structured surgical video data can lead to severely underestimated model uncertainty. To ensure relevance across both low-level and high-level prediction tasks, we focused on two widely used benchmark tasks, instrument segmentation (RobustMIS) and surgical action recognition (CholecT45). Both datasets exhibit a clear hierarchical structure, with multiple correlated frames per patient case.
For the binary segmentation task, we used the results of the ten algorithms of the binary segmentation task of the RobustMIS challenge. The same metrics as in the original challenge were applied, namely the DSC and NSD. In this dataset, one hierarchical level was considered, namely the patient (i.e., the video; n = 10).
For the surgical action triplet recognition task, results were derived from the CholecT45 dataset using one algorithm (Swin-Base Transformer; see above). We calculated the original mAP [maier2024metrics] metric as done in the original challenge. Given the class imbalance across 100 triplet classes, we additionally computed a class-weighted mAP, and included the top-5 Accuracy for completeness. Metrics were calculated for each cross-validation fold. Again, we considered the patient (i.e., the video; n = 45) as the relevant level of hierarchy.
For both tasks, CIs were estimated using two resampling strategies: the standard bootstrap and the hierarchical bootstrap. In the standard (naïve) bootstrap, we performed resampling with replacement of all frames of the entire test set 1,000 times without considering the hierarchical structure [efron1994introduction]. For each resample, we calculated the mean metric for each bootstrap sample, and obtained the empirical quantiles from the resulting bootstrap distribution to calculate 95% CIs. For the surgical action triplet recognition task, resampling was applied across all frame-level predictions within each cross-validation fold.
In contrast, the hierarchical bootstrap explicitly accounted for dependencies at each hierarchy level [saravanan2020application]. We first resampled the videos, i.e., the higher hierarchy level, followed by resampling the individual frames within each sampled video. The mean metric was then computed across all resampled frames and videos. This process was repeated 1,000 times, and the empirical quantiles of the resulting metric means were used to estimate the CIs. For the surgical triplet recognition task, this procedure was applied separately within each cross-validation fold, with metrics averaged per video and per fold before CI estimation.
@@
This experiment investigated whether global (non-stratified) aggregation of metric scores can conceal algorithm weaknesses under challenging image conditions. To enable stratified analysis across clinically relevant image characteristics, we focused on the RobustMIS dataset, for which we had access to structured metadata on visual artifacts or image properties [ross2023beyond]. While the original study [ross2023beyond] employed these annotations to analyze model robustness across visual conditions, our analysis focused on how global aggregation can obscure property-dependent performance differences that are critical for assessing validation reliability. Specifically, we analyzed the multi-instance segmentation task of the RobustMIS challenge, using the MI_DSC scores from the seven participating algorithms.
For each algorithm, the median MI_DSC across all frames was calculated as the baseline performance (non-stratified). The median was chosen instead of the mean to reduce sensitivity to outliers. Stratified performance was then calculated by restricting the analysis to frames containing individual artifact types. The following properties were considered: blood, reflections, smoke, motion, overexposed, underexposed, intersecting instruments, and low-artifact scenes (i.e., scenes with one or fewer annotated properties). Because frames could contain multiple artifacts simultaneously, these subsets were not mutually exclusive.

### v2

@@ 3. Discussion
Our work provides a comprehensive, expert-driven taxonomy of validation pitfalls in surgical AI, supported by empirical evidence and experimental results. By linking methodological flaws to surgical risks, our framework highlights the need for rigorous validation to ensure safe and effective AI deployment. Our findings demonstrate that common validation practices frequently ignore the temporal and hierarchical structure of surgical data, leading to overconfident or clinically irrelevant conclusions. Through a Delphi process with experts from surgery, machine learning, biostatistics, and regulatory affairs, we specifically identified and contextualized 18 critical pitfalls. This multi-stakeholder approach ensured that the collected pitfalls are both technically sound and clinically meaningful, making them actionable for a broad range of players, from algorithm developers to regulatory reviewers.The added value of the present work is four-fold:
We provide a comprehensive and structured single point of access to validation pitfalls in surgical AI, making existing and newly contextualized insights accessible to surgical data scientists and clinicians.
Beyond conceptual discussion, we quantify the frequency and distribution of these pitfalls specifically within the SDS literature, thereby providing empirical evidence on how common these issues are in current practice.
We empirically demonstrate the practical relevance of selected pitfalls through experiments on real surgical data, showing how inappropriate validation choices can substantially affect reported performance, uncertainty estimates, and algorithm rankings.
We provide practical, pitfall-specific best practices, translating the identified shortcomings into actionable guidance for improved validation practice.
It should be noted that while some of the identified validation pitfalls have previously been discussed in the broader computer vision and machine learning literature, these insights are typically scattered across domain-agnostic, technically dense publications and are not tailored to the specific characteristics of surgical data science. Unlike prior work that discussed individual pitfalls in isolation, our study combines expert consensus, empirical literature screening, and experimental analysis to provide a structured, systematic, and evidence-based single point of access for validation pitfalls in surgical video AI.
@@
The collected pitfalls span all stages of the validation process. For example, at the data level, geographical imbalance of surgical and endoscopic datasets (Extended Data Fig. 1b) shows how limited geographic diversity can restrict representativeness, as many public benchmarks have historically been acquired from a narrow range of regions. Recent initiatives, such as the Critical View of Safety (CVS) Challenges 2024 and 2025 [Alapatt et al., 2025], represent encouraging steps toward broader global inclusion, yet most benchmark datasets still predominantly reflect surgical practice patterns from Western regions. Beyond geographical and institutional biases, non-representativeness in surgical video data may also arise from unobserved operator-related heterogeneity. In particular, videos recorded in academic or training centers may involve multiple operators with different experience levels (e.g., trainees and attending surgeons) within a single procedure, introducing additional variance if operator roles are not annotated or reported. At the metric level, misalignment between chosen metrics and clinical objectives remains common, as metrics tailored to surgeon-specific needs or temporal aspects are often lacking. Consequently, studies frequently rely on simple frame-based scores such as Accuracy or DSC, which fail to capture the temporal and hierarchical complexity of surgical workflows and can obscure clinically relevant weaknesses. At the aggregation level, naïve frame-wise aggregation across temporally dependent data can mislead performance estimates, concealing critical failure modes and even altering algorithm rankings.
@@
In addition to establishing this taxonomy, we collected empirical evidence confirming that the identified pitfalls are widespread across the surgical AI literature. Our systematic screening revealed frequent issues such as unstratified aggregation, lack of uncertainty reporting, unjustified and unsuitable metric selection, and poor documentation of validation procedures. Despite temporality being inherent to intraoperative surgical workflows and despite the existence of temporal metrics, these were rarely used in practice. Generally speaking, although surgical video AI tasks often involve complex, structured data such as multiple classes, temporal dependencies, and hierarchical structure, validation in common practice frequently remains limited to simple Accuracy metrics. As shown by our experiments, naïve frame-wise validation without consideration of temporal aspects or data hierarchy can mask critical failure modes and lead to a substantial underestimation of uncertainty. Together, these findings underscore the need for surgical AI.
Building on our findings, we derived practical, pitfall-specific best practices through expert consensus in the Delphi process. The best practice recommendations encompass measures to avoid specific pitfalls and strategies to address them transparently when they cannot be fully avoided. Their applicability depends on the study context and constraints, such as available data and metadata, annotation resources, and intended use. Accordingly, the scope and relevance of individual best practices vary across stages of method development, ranging from early exploratory studies to benchmarking and evaluations supporting claims of generalizability or clinical readiness. Our collection is intended to support informed selection of best practices that are meaningful for a given use case rather than an exhaustive checklist or compliance framework, while encouraging transparent documentation of remaining limitations when constraints prevent full implementation. The suggested best practices could serve as a useful resource for writing and reviewing scientific articles in the domain of SDS.
Beyond technical adherence to best practices, the clinical translation of surgical AI requires alignment with evolving regulatory frameworks and the multifaceted needs of healthcare stakeholders. A critical challenge in this transition is the fundamental difference between diagnostic AI and interventional AI. While diagnostic AI often produces a single, static prediction (e.g., ”lesion present”), interventional AI operates within a dynamic, time-dependent environment. In intraoperative applications in surgery, the causal link between a specific model output at a single timepoint (e.g., a ”cut” vs. ”no-cut” recommendation) and a subsequent clinical outcome or complication ares often difficult to isolate. This fuzzy relationship in error attribution poses hurdles for traditional validation paradigms and regulatory bodies, which typically seek a direct mapping of model performance to patient safety. Consequently, satisfying stakeholder requirements for clinical readiness will necessitate validation strategies that move beyond aggregate metrics toward implementation science approaches. These must account for the temporal dynamics of the surgical workflow and the complex interaction between the AI’s guidance and the surgeon’s intraoperative decision-making, ensuring that validation mirrors the high-stakes, real-time nature of the operating room.
@@
Our work comes with several limitations. While our taxonomy targets surgical video analysis, its generalizability to other temporally structured domains (e.g., cardiology) remains to be explored. Methodologically, although our Delphi expert consortium included more than 90 international experts and covered a broad range of expertises, it may not have captured the full diversity of surgical subfields and regulatory perspectives. Participation also varied across Delphi rounds, raising the potential for biases in the weighting and selection of pitfalls. In addition, our systematic literature review covered only 46 articles. However, although the sample size was limited, MICCAI is the leading conference for medical image analysis and computer-assisted interventions and can thus be considered representative for the field. To mitigate this limitation and assess whether our findings extend beyond a single conference year, we performed an additional targeted screening of 89 MICCAI 2025 SDS papers. This analysis was conducted by a single rater and focused on the key pitfalls reported in the Results rather than the full set of parameters. While less comprehensive than our primary screening, it yielded highly similar trends across the evaluated aspects, supporting the robustness of our main findings. It should further be noted that our experiments only tackled selected pitfalls. While pitfalls such as data leakage have been demonstrated in the broader machine learning community, we explicitly focused on aspects that are especially critical in surgery: aggregation under temporal and hierarchical structures. Other pitfalls are illustrated in smaller real-data analyses in Extended Data Figs. 1- 9. Finally, our present work does not yet provide concrete solutions for each pitfall. However, developing concrete recommendations will be the focus of future work of the consortium.
@@
Several open research directions emerge from our presented pitfall taxonomy. At a conceptual level, future work should move beyond surrogate metrics toward validation that reflects clinical benefit and patient outcomes. This includes defining what constitutes sufficient performance within specific clinical contexts and establishing comparability across heterogeneous tasks. Achieving this goal requires close collaboration between AI developers and clinical experts. In the surgical domain, only such multidisciplinary input enables a meaningful interpretation of validation metrics in relation to workflow impact, safety considerations, and patient outcomes. Methodologically, clear validation phases for SDS systems that integrate governance and stakeholder input, as well as standardized reporting, need to be defined and appropriately integrated into clinical trial design. Ensuring robustness through post-deployment monitoring, addressing catastrophic-failure risk and safety considerations, and enabling effective human-AI collaboration in the OR will be equally crucial. From a technical and adoption perspective, progress depends on harmonizing label ontologies and annotation protocols across datasets, facilitating validation for multimodal data while protecting patient privacy, and assessing behavioral consistency of AI models across samples and software versions to ensure stability after updates or retraining. Finally, embedding clinician priorities, workflow impact, and real-time safety mechanisms should become integral validation goals.
@@
In summary, while Metrics Reloaded [Maier-Hein et al., 2024] provided metric recommendations for image-based validation, our new framework extends this foundation to pitfalls stemming from the temporal and hierarchical complexity of surgical video analysis. We envision this work as a catalyst for improved validation practice and future benchmarking efforts. By raising awareness of widespread pitfalls, we aim to inspire a paradigm shift toward more robust, interpretable, and clinically grounded validation pipelines. By systematically mapping validation pitfalls to their consequences and by providing practical best practice recommendations, this work can serve as a structured foundation for integrating validation quality criteria into clinical trial design, regulatory review, and publication guidelines. Concretely, our findings may inform the development or refinement of reporting standards for medical video analysis, such as TRIPOD-AI [Collins et al., 2021], DECIDE-AI [Vasey et al., 2022], or future domain-specific extensions. Going forward, our research consortium will focus on translating these pitfalls into surgery-specific metric and aggregation recommendations, further advancing the reliability and clinical readiness of surgical AI models.
@@ 4. Methods
4.1. Identification of validation pitfalls and best practices through a multi-stage Delphi process and complementary searches
@@
The pitfalls presented in this work were derived through a combination of approaches, centered on a multi-stage, consensus-driven Delphi process conducted by an international, multidisciplinary panel of experts. A Delphi process is a structured consensus-building approach in which experts provide input individually – typically through questionnaires – followed by rounds of controlled feedback and refinement [Brown, 1968]. This methodology is widely recognized in medicine as a way to establish best practices in areas in which the available evidence is limited, inconsistent, or missing [Nasa et al., 2021].
@@
Our Delphi panel initially included 60 international experts from the SDS initiative. To broaden the diversity of the expertise, the consortium was gradually expanded to 92 members across 68 institutions, reflecting both technical and clinical backgrounds. The expert panel was composed of 30% clinical, 75% technical, and 5% shared expertise. 12% of experts were from industry. The majority of experts were affiliated in Europe (70%; mostly Germany (36%) and United Kingdom (12%)) and North America (25%; mostly United States of America (20%)), followed by Asia (7%) and Africa (2%).
@@
This initiative started in March 2023 with an initial scoping survey to identify the most critical problems in validating surgical AI and to capture use cases lacking suitable metrics or showing discrepancies between metrics and the clinical needs (participation rate: 33%). Building on the survey findings, we held an in-person kickoff workshop at the SAGES annual meeting in Montréal, Canada (41% in-person participation). This workshop refined the project scope and set priorities, laying the groundwork for the following Delphi rounds. Based on the workshop discussions, participants agreed to focus the initiative on surgical video understanding, reflecting shared priorities across clinical and technical stakeholders. Following the workshop, the core team performed targeted literature searches to compile candidate pitfalls. In parallel, a joint retreat involving members from three research groups at the German Cancer Research Center (DFKZ), the National Center for Tumor Diseases (NCT), and the University College London (UCL) provided an additional forum to critically discuss and refine preliminary pitfalls based on practical experience and interdisciplinary perspectives.
@@
In addition, the core team performed a literature review to identify additional pitfalls. Concretely, we explored two databases, namely PubMed and GoogleScholar, as well as a general Google search, and used the following search string: (”surgical data science” OR ”surgical artificial intelligence” OR ”surgical AI” OR ”surgical scene understanding” OR ”surgical video analysis”) AND (”validation” OR ”evaluation” OR ”metric”) AND (”pitfall” OR ”limitation” OR ”caveat” OR ”drawback” OR ”shortcoming” OR ”weakness” OR ”flaw” OR ”disadvantage”). PubMed returned no relevant results, while Google Scholar yielded 704 hits and a general Google search 94 results (30 non-peer-reviewed). Although these searches yielded several appropriate items, no comprehensive or structured collection of validation pitfalls was identified.
@@
In total, we conducted six Delphi rounds (participation rates: 64%, 70%, 48%, 53%, 46%, 60%). Round 1 confirmed the overall project scope, while round 2 refined the pitfall categorization and pitfall list, identified missing pitfalls, and collected supporting references for the evidence of pitfalls. Round 3 focused on linking pitfalls to consequences and risks, as well as an initial collection of best practices, and round 4 sought final consensus on the pitfall catalog and pitfall categorization (agreement: 98%), as well as optional feedback on figures, experiments (agreement 90%), and open research questions. The core team refined the best practices collected from round 3, which were finalized in rounds 5 and 6, where only items with an agreement above 80% were kept. In the final round, agreement for the list of best practices was above 90% for each pitfall.
@@ 4.2. Systematic review for prevalence of pitfalls
From all MICCAI 2023 papers (n = 730), we identified those papers related to SDS (n = 51). From those, five articles were excluded because these did not deal with deep learning-related methods, therefore, several questions did not apply (remaining n = 46). Each paper was screened by two independent screeners, from a total of twelve screeners. Afterwards, a third senior screener compared the results and resolved conflicts. In total, three senior screeners joined in this last step, the papers were divided among them. The screening covered more general aspects such as task or surgery type, but specifically focused on evidence for the identified pitfalls. In line with the Delphi process and after identifying the final list of pitfalls, a follow-up screening, following the same process, was conducted to ensure evidence for all pitfalls.
To assess the robustness of our findings over time, we additionally performed a targeted screening of MICCAI 2025 SDS papers (n = 89). In contrast to the primary analysis, this screening was conducted by a single rater and focused on the subset of pitfalls reported in the Results section. The purpose was not to repeat the full screening procedure, but to evaluate whether the main findings hold across a more recent set of publications.
@@ 4.3. Experimental Design
To ensure the general relevance of our findings, we based our experiments on two widely used datasets in surgical video analysis [Carstens et al., 2025] that have been used in international challenges, that are highly cited in the field, and cover two key tasks in surgical video analysis: instrument segmentation and action recognition.
The RobustMIS challenge 2019 [Roß et al., 2021] consists of videos from 30 laparoscopic colorectal surgeries, covering three surgery types, namely rectal resection, proctocolectomy, and sigmoid resection, with 10 videos, i.e., patients, per surgery type. For the challenge, the data from rectal resection and proctocolectomy surgeries were used for training and internal testing (stages 1 and 2), while sigmoid resection cases (stage 3) were reserved for assessing generalization to an unseen surgery type. For our experiments, we restricted the analysis to data from stage 3 (sigmoid resection). The challenge consisted of three tasks: binary segmentation, multi-instance detection, and multi-instance segmentation. For our experiments, we leveraged results for both segmentation tasks. For binary segmentation, a total of ten algorithms were involved, for multi-instance segmentation, seven algorithms participated. The challenge metrics included the DSC, which measures the overlap between prediction and reference, and NSD, which assesses boundary accuracy [Maier-Hein et al., 2024] for binary segmentation and their multi-instance variants (MI_DSC and MI_NSD) for the multi-instance segmentation task. We had access to the frame-level metric scores submitted by all participating teams, which allowed us to analyze the impact of validation choices across a diverse set of real-world algorithms. All results were used in anonymized form to ensure confidentiality.
In addition to the segmentation masks, we utilized additional structured meta-annotations for the RobustMIS data, describing the presence of common visual artifacts or image characteristics [Roß et al., 2023]. These meta annotations indicate whether specific challenges such as blood, smoke, or motion blur are present per frame and instrument.
We further used the data and task setup of the CholecTriplet challenge [Nwoye et al., 2023a], which is based on 45 laparoscopic cholecystectomy videos (CholecT45). The task involves recognition of surgical action triplets, with annotations for 100 triplet classes, each defined by a combination of instrument, verb, and target. For our experiments, we used a Swin-Base Transformer trained using multitask learning, incorporating information on the instrument, verb, and target and soft-labels generated using a multi-teacher approach [Yamlahi et al., 2023]. Model validation was performed using 5-fold cross-validation, following the official CholecT45 setup.
@@
This experiment investigated how ignoring data dependencies in temporally structured surgical video data can lead to substantially underestimated model uncertainty. To ensure relevance across both low-level and high-level prediction tasks, we focused on two widely used benchmark tasks, instrument segmentation (RobustMIS) and surgical action recognition (CholecT45). Both datasets exhibit a clear hierarchical structure, with multiple correlated frames per patient case.
For the binary segmentation task, we used the results of the ten algorithms of the binary segmentation task of the RobustMIS challenge. The same metrics as in the original challenge were applied, namely the DSC and the NSD. In this dataset, one hierarchical level was considered, namely the patient (i.e., the video; n = 10).
For the surgical action triplet recognition task, results were derived from the CholecT45 dataset using one algorithm (Swin-Base Transformer; see above). We calculated the original mAP [Maier-Hein et al., 2024] metric as done in the original challenge. Given the class imbalance across 100 triplet classes, we additionally computed a class-weighted mAP, and included the top-5 Accuracy for completeness. Metrics were calculated for each cross-validation fold. Again, we considered the patient (i.e., the video; n = 45) as the relevant level of hierarchy.
For both tasks, CIs were estimated using two resampling strategies: the standard bootstrap and the hierarchical bootstrap. In the standard (naïve) bootstrap, we performed resampling with replacement of all frames of the entire test set 1,000 times without considering the hierarchical structure [Efron and Tibshirani, 1994]. For each resample, we calculated the mean metric for each bootstrap sample, and obtained the empirical quantiles from the resulting bootstrap distribution to calculate 95% CIs. For the surgical action triplet recognition task, resampling was applied across all frame-level predictions within each cross-validation fold.
In contrast, the hierarchical bootstrap explicitly accounted for dependencies at each hierarchy level [Saravanan et al., 2020]. We first resampled the videos, i.e., the higher hierarchy level, followed by resampling the individual frames within each sampled video. The mean metric was then computed across all resampled frames and videos. This process was repeated 1,000 times, and the empirical quantiles of the resulting metric means were used to estimate the CIs. For the surgical triplet recognition task, this procedure was applied separately within each cross-validation fold, with metrics averaged per video and per fold before CI estimation.
@@
This experiment investigated whether global (non-stratified) aggregation of metric scores can conceal algorithm weaknesses under challenging image conditions. To enable stratified analysis across clinically relevant image characteristics, we focused on the RobustMIS dataset, for which we had access to structured metadata on visual artifacts or image properties [Roß et al., 2023]. While the original study [Roß et al., 2023] employed these annotations to analyze model robustness across visual conditions, our analysis focused on how global aggregation can obscure property-dependent performance differences that are critical for assessing validation reliability. Specifically, we analyzed the multi-instance segmentation task of the RobustMIS challenge, using the MI_DSC scores from the seven participating algorithms.
For each algorithm, the median MI_DSC across all frames was calculated as the baseline performance (non-stratified). The median was chosen instead of the mean to reduce sensitivity to outliers. Stratified performance was then calculated by restricting the analysis to frames containing individual artifact types. The following properties were considered: blood, reflections, smoke, motion, overexposed, underexposed, intersecting instruments, and low-artifact scenes (i.e., scenes with one or only fewer annotated properties). Because frames could contain multiple artifacts simultaneously, these subsets were not mutually exclusive.